mshake:

Explain how you would graph the line containing a slope of –1/5 that goes through the point (1,–4).

Mercury:

Please do not copy-paste my response. It is stated that the graph goes through the point (1,–4), so you could start by plotting that point. Now, for the slope (-1/5), remember that slope = rise/run. slope (-1/5) implies rise of -1 (so down 1 unit) and run of 5 (so to the right 5 units).* so start with (1,-4) which has just been plotted, go down 1 unit, then across 5 units, and draw the new point there. from there, you can connect the two points with a straight line.

Mercury:

* note: slope (-1/5) could also imply rise = 1 and run = -5, which will go in the opposite direction but still generate the same slope. so this could also work to generate another point on the graph.